# Hardness-Dependent Optimized Quantum Adiabatic Schedules for Maximum Independent Sets

### v1

@@ Abstract
We propose a numerical approach to design highly efficient schedules for adiabatic quantum computing, focusing on the maximum independent set problem and on neutral atom platforms. Based on a representative dataset of small graphs, we present numerical evidences that the optimum schedules depend principally on the hardness of the problem and not on its size. These schedules perform better than the benchmark protocols and admit a straightforward implementation in the hardware. This allows us to extrapolate the results to larger graphs and to successfully solve moderately hard instances with more than 100 atoms using QuEra’s quantum computer “Aquila”. We believe that extending our approach to hybrid algorithms could be the key to solve the hardest problems with the current technology.
maximum independent set; neutral atoms; QuEra’s Aquila; analog quantum computing
I Introduction
@@
In recent years, quantum computers have emerged as a promising way to solve complex optimization tasks in science and industry [1, 2, 3, 4]. Typically, the problem of interest is encoded in the ground state of a quantum Hamiltonian that can be reached through adiabatic evolution, starting from an easy-to-prepare initial state [5, 6, 7, 8]. Advanced protocols have to be developed to overcome not only the experimental limitations, such as qubit decoherence and control errors, but also the intrinsic complication of vanishing gap in the eigenspectrum during the evolution [9, 10, 11, 12].
@@
Here, we focus on the maximum independent set (MIS) problem, where the task is to find the largest set of vertices that do not share an edge in a graph. Such problems appear naturally in practical applications, for instance while designing resilient communication networks [13], and are known to be NP-hard in general [14]. Finding a MIS is still a difficult task in the particular case of unit-disk graphs, in which vertices are connected by an edge if they are closer than a fixed distance [15]. This family of graphs admits a direct mapping to arrays of Rydberg atoms, which explains the popularity of the MIS problem in analog quantum computing [16, 17, 18].
@@
In this setting, optical tweezers enable arbitrary arrangements of atoms and controlled laser pulses manipulate the quantum states [19, 20]. These lasers drive the adiabatic evolution of the system following schedules that are specific to the MIS problem. The precise description of the schedules is of utmost importance since it conditions the success of the quantum algorithm. Many methods have been proposed in that regard: quantum approximate optimization algorithm and variational quantum eigensolver [21, 22, 23, 24], shortcut to adiabadicity and counterdiabatic protocol [25, 26, 27], Bayesian optimization and machine learning [28, 29], to name a few.
@@
Most aforementioned methods require multiple iterations to be run on a quantum device and to be updated by a classical optimizer and thus belong to the class of hybrid algorithms [30]. By their very nature, these algorithms cost both time and money as they make intense use of the quantum hardware. In this study, we show how classical precomputation of optimized schedules for small graphs can lead to highly efficient protocols for large graphs on the quantum device. First, we introduce the mathematical concepts that are used throughout this work. Then, in Sec. II we describe how to optimize the schedules based on numerical simulations of small graphs, and in Sec. III we present the results obtained in experiments. Detailed calculations and additional results are presented in the appendix.
@@
The analog quantum dynamics is governed by the Hamiltonian $H=H_{\text{drive}}+H_{\text{cost}}$, where the driving part and the cost function are given by [18]
@@
In these equations, $\Omega$ is the Rabi drive amplitude, $\phi$ its phase, $\Delta$ the detuning, and $\hat{n}_{i}=\ket{1}_{i}\bra{1}$ | detects the Rydberg excitation of the atom $i=1,\ldots,N$. The two-body potential is $V_{ij}=C_{6}/|\textbf{r}_{i}-\textbf{r}_{j}|^{6}$, with $C_{6}$ a constant for the Van der Waals interaction between two Rydberg states at positions $\textbf{r}_{i}$ and $\textbf{r}_{j}$, which encodes the unit-disk graph of the MIS problem to be solved. The boundary conditions at initial and final times $t_{\text{i}}$ and $t_{\text{f}}$ for an adiabatic evolution are usually stated as follows:
@@
In practice, we implement the first inequality as $\Delta_{\text{i}}\leq-\delta_{\Delta}$, where $\delta_{\Delta}$ is a small positive constant that incorporates the main sources of error in the quantum hardware, so that all atoms are ensured to start in the ground state |$\ket{0}$. More importantly, we show in Sec. SI that the final detuning should lie in a narrow interval $[\Delta_{\text{LB}},\Delta_{\text{UB}}]$, where the bounds are explicitly given by
$\frac{C_{6}}{12\,a^{6}}\leq\Delta_{\text{f}}\leq\frac{C_{6}}{8\,a^{6}},$ (3)
with $a$ the lattice spacing. Note that these bounds are not universal, but they are satisfied by the vast majority of unit-disk graphs on a square lattice.
@@
A detailed analysis of the dynamics of the (classical) simulated annealing for MIS reveals that the graphs with many suboptimal solutions are likely to trap the algorithm in local minima, hence missing the global solution [18]. This observation leads to the following definition of the hardness parameter of a graph:
@@
II Numerical optimization
In this section, we consider a fixed set of 500 representative graphs of size 8 to 17  [^1] We equally refer to the number of vertices in a graph as order or size. In quantum computing, the latter is commonly understood as the size of the system and not as the number of edges in the graph (mathematical definition)., which allows us to perform fast classical simulations and to use efficient deterministic optimization algorithms. This dataset is generated in such a way that the correlation between the size of the graphs and their hardness parameter is minimized, see Sec. SIII. Therefore, the results that depend principally on $\mathcal{HP}$ may be extrapolated, in principle, to graphs of any size. We focus on the probability to reach a MIS solution at the end of the adiabatic evolution, which we denote by $P_{\text{MIS}}$ and use as the score function to be optimized. Note that most studies consider the approximation ratio instead (normalized mean energy of the final state), but in our opinion only the lowest energy level and not the full energy spectrum has a practical utility because of the very definition of the hardness parameter, see Eq. (4). More details on the dataset and the numerical methods are given in Sec. SIII.
Following the “best practices” listed in [20], we consider protocols for which the Rabi drive $\Omega(t)$ reaches $\Omega_{\max}$ during the evolution, so that the associated dynamic blockade radius becomes $R_{\text{b}}=(C_{6}/\Omega_{\max})^{1/6}\approx 8.367$. In order to generate a unit-disk graph on nearest and next-nearest neighbors, this radius should satisfy $\sqrt{2}a<R_{\text{b}}<2a$, which in turn implies $5.9164.183<a<$ 5.916. We set $a=5$ which yields good results in preliminary calculations. We also consider a fixed evolution time $t_{\text{f}}=1$ so that $P_{\text{MIS}}$ varies as much as possible depending on the various protocols; much smaller or larger times would lead to either extremely bad or extremely good results independently of the precise shape of the schedules.
@@ II Numerical optimization
II.1 Piecewise linear protocol
The typical adiabatic protocol is based on piecewise linear schedules, where $\Omega$ is slowly ramped up, then $\Delta$ is ramped from negative to positive, and finally $\Omega$ is ramped down, see [20, Fig. 3.1b] or [34, Fig. 2c]. Usually, these schedules are symmetric in the sense that the initial and final ramp times $\tau_{\text{i}}$ and $\tau_{\text{f}}$ are equal, and that the initial and final detunings $\Delta_{\text{i}}$ and $\Delta_{\text{f}}$ have opposite values. In our case, we are looking for schedules that are as general as possible, so we keep all parameters independent. Moreover, we consider distinct values of the ramp times for $\Omega$ and $\Delta$. After few optimization steps, it appears that the ramp times of the detuning tend to vanish; we therefore set them to zero. Then, we optimize the four remaining parameters $(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for each graph of the dataset. One interesting result is that the optimum parameters mostly depend on the hardness parameter and not on the order of the graph, see Fig. 1. The consequence is twofold: first, we can fit these data to create a simple model of optimized schedules, se Fig. 2, and second, the results obtained from small graphs can be extrapolated to larger graphs (Sec. III.2). In other words, many costly and time consuming calls to quantum computers that are required in hybrid optimization may be saved by using parameters that are designed by purely classical means.
@@ II.1 Piecewise linear protocol
Figure 2: Optimized piecewise linear schedules. These functions depend on the hardness of the graph and are illustrated here for $\mathcal{HP}\in\{0.5,1,2,4,8\}$ (thin lines, from green to maroon). Any large graph of unknown hardness is best solved by using the parameters of the limiting case $\mathcal{HP}\to\infty$ (dark maroon thick line): $\tau_{\text{i}}\approx t_{\text{f}}/2$, $\tau_{\text{f}}\approx t_{\text{f}}/10$, $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/2$, $\Delta_{\text{f}}=\Delta_{\text{UB}}$.
@@
The aim of the adiabatic evolution is to find the MIS, which only then allows one to calculate the value of $\mathcal{HP}$. In the case of small graphs, since all vertex configurations are enumerated easily, we consider that $\mathcal{HP}$ is known and that it can be used to choose the best schedules based on the model. For large graphs, however, the hardness parameter cannot be considered to be a known value. In that case, one should either use the schedules with $\mathcal{HP}\to\infty$, which yields quite good results for all graphs, or eventually consider a one-dimensional search on the hardness parameter, which is still much easier than performing a multi-dimensional optimization of all schedule parameters.
Further improvement
The linear schedules depicted in Fig. 1 yield very good results compared to existing protocols, which motivates us to go one step further by considering not 4 but 6 free parameters, see Sec. SIII.3. The optimized schedules are still easily implementable, yet they perform much better than all benchmark protocols, see Sec. II.3.
II.2 Couterdiabatic protocol
The limited coherence time of the quantum state in the quantum hardware is a major source of error in adiabatic computing. Shortcuts to adiabaticity are well-known methods to overcome such limitations [35]. Among such methods, the counterdiabatic (CD) driving has gained considerable attention over the last years [36, 25, 37]. The idea of the CD driving is to add a gauge potential $\mathcal{A}$ to the Hamiltonian $H$ such that the transitions between eigenstates are suppressed:
$H_{\text{CD}}(t)=H(t)+\mathcal{A}(t).$ (5)
@@ II.2 Couterdiabatic protocol
The calculation of the exact gauge potential is a very difficult task in general, and these mathematical terms may not be always adapted to an experimental implementation. Neglecting the interaction terms in the calculations, one can however find a suitable gauge potential which utilizes the phase of the Rabi drive [27]. We generalize the smooth schedules of this so-called Analog Counterdiabatic Quantum Computing (ACQC) protocol as follows:
$\Omega(t)$ $=\Omega_{\max}\,\sin\left(\frac{\pi}{2}\,\sin\left(\theta_{\Omega}(t)\right)\right)^{2},$ (6a)
$\Delta(t)$ $=\frac{\Delta_{\text{i}}+\Delta_{\text{f}}}{2}+\frac{\Delta_{\text{i}}-\Delta_{\text{f}}}{2}\cos\left(\theta_{\Delta}(t)\right),$ (6b)
where $\theta(t)$ is any monotonic smooth function that ranges from 00 to $\pi$; here we set $\theta_{\Omega,\Delta}(t)=\pi\,t/t_{\text{f}}$. Moreover, based on the variational procedure proposed in [26], we show in Sec. SII how to improve the standard ACQC protocol to explicitly take into account the graph-specific coefficients $V_{ij}$. We consider an additional free parameter $0\leq\nu\leq C_{6}/a^{6}$ that interpolates between these two gauge potentials, which greatly helps in the optimization process.
As for the piecewise linear protocol, we optimize the schedule parameters for all graphs of the dataset. We observe the same dependence on the hardness parameter (Fig. 3) as well as a strong shift of the initial and final detunings (Fig. S7). We also tried to improve the results by considering more evolved $\theta$ functions (in particular, to mimic the skewed shape of the Rabi amplitude in Fig. 2), but no significant difference was found in this case.
@@
In Aquila, the Rabi amplitude and the detuning must be defined as piecewise linear and not smooth functions, while the phase must be piecewise constant. However, the step size can be as small as 0.05 µs, which results in maximum 80 intervals for a coherent evolution. The schedules are therefore well approximated in the quantum hardware, but one could slightly improve the experimental results by taking this hardware capability into account in the numerical optimization.
@@
Figure 3: Optimum parameters for the CD schedules. The initial detuning (top) reaches the limit $-\delta_{\Delta}$ for several graphs, but in average it tends to $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/10$ for the hardest graphs. The final detuning (bottom) also increases with the hardness parameter and quickly reaches the limit $\Delta_{\text{UB}}$. Note that the easy graphs exhibit optimized values of $\Delta_{\text{f}}$ that are smaller than the statistically determined constant $\Delta_{\text{LB}}$. This is due to the fact that the exact graph-dependent lower bounds are much smaller in that case, see Fig. S2. Finally, the interpolating coefficient (not shown) displays a similar behavior and tends to a strictly positive value $\nu=4.30$ in the current setting, indicating that the generalized gauge potential plays an important role in this MIS protocol.
II.3 Results

### v2

# Maximum Independent Set; Hardness Parameter; Neutral Atoms; Rydberg Blockade; QuEra’s Aquila; Analog Quantum Computing; Adiabatic Schedules; Counterdiabatic Driving

@@ Abstract
We propose a numerical approach to design highly efficient adiabatic schedules for analog quantum computing, focusing on the maximum independent set problem and neutral atom platforms. Based on a representative dataset of small graphs, we present numerical evidences that the optimum schedules depend principally on the hardness of the problem rather than on its size. These schedules perform better than the benchmark protocols and admit a straightforward implementation in the hardware. This allows us to extrapolate the results to larger graphs and to successfully solve moderately hard problems using QuEra’s 256-qubit Aquila computer. We believe that extending our approach to hybrid algorithms could be the key to solve the hardest instances with the current technology, making yet another step toward real-world applications.
1 Introduction
@@
In recent years, quantum computers have emerged as a promising way to solve complex optimization tasks in science and industry [1, 2, 3, 4]. Typically, the problem of interest is encoded in the ground state of a quantum Hamiltonian, which can be reached through adiabatic evolution starting from an easy-to-prepare initial state [5, 6, 7, 8]. Advanced protocols have to be developed to overcome not only the experimental limitations, such as qubit decoherence and control errors, but also the intrinsic complication of vanishing gap in the eigenspectrum during the evolution [9, 10, 11, 12].
@@
Here, we focus on the maximum independent set (MIS) problem, where the task is to find the largest set of vertices that do not share an edge in a graph. Such problems appear naturally in practical applications, for instance while designing resilient communication networks [13], and are known to be NP-hard in general [14]. Finding a MIS is still a difficult task in the particular case of unit disk graphs [15, 16], in which vertices are connected by an edge if they are closer than a fixed distance. This family of graphs admits a direct mapping to arrays of Rydberg atoms, which explains the popularity of the MIS problem in analog quantum computing [17, 18, 19]. In that setting, optical tweezers enable arbitrary arrangements of atoms and controlled laser pulses manipulate the quantum states [20, 21]. The lasers then drive the evolution of the system according to schedules that are specific to the problem to be solved.
@@
The precise description of the adiabatic schedules is of utmost importance since it conditions the success of the quantum algorithm. Many methods have been proposed in this regard: quantum approximate optimization algorithm and variational quantum eigensolver [22, 23, 24], shortcut to adiabadicity and counterdiabatic protocol [25, 26, 27, 28], Bayesian optimization and machine learning [29, 30], to name a few. Most aforementioned methods require multiple iterations to be run on a quantum device and to be updated by a classical optimizer and thus belong to the class of hybrid algorithms [31]. By their very nature, these algorithms cost both time and money as they make intense use of the quantum hardware.
@@
In this study, we show how a precomputation of optimized schedules for small graphs can lead to highly efficient protocols for graphs of larger size on the quantum device [^1] We equally refer to the number of vertices in a graph as order or size. In quantum computing, the latter is commonly understood as the size of the system and not as the number of edges in the graph (mathematical definition).. First, we introduce the basic concepts that are used throughout this work. Then, in Sec. 2 we describe how to optimize the schedules based on numerical simulations of small graphs, and in Sec. 3 we present the results obtained in experiments. Detailed information and calculations are available in the appendix.
@@
The analog quantum dynamics is governed by the Hamiltonian $H=H_{\text{drive}}+H_{\text{cost}}$, where the driving part and the cost function are [19]
@@
In these equations, $\Omega$ is the Rabi drive amplitude, $\phi$ its phase, $\Delta$ the detuning, and $\hat{n}_{i}=\ket{1}_{i}\bra{1}$ | detects the Rydberg excitation of the atom $i=1,\ldots,N$. The two-body potential is $V_{ij}=C_{6}/|\textbf{r}_{i}-\textbf{r}_{j}|^{6}$, with $C_{6}$ a constant for the Van der Waals interaction between two Rydberg states at positions $\textbf{r}_{i}$ and $\textbf{r}_{j}$, encoding the unit disk graph of the MIS problem. The boundary conditions at initial and final times $t_{\text{i}}$ and $t_{\text{f}}$ for the adiabatic evolution are usually stated as follows:
@@
In practice, we implement the first inequality as $\Delta_{\text{i}}\leq-\delta_{\Delta}$, where $\delta_{\Delta}$ is a small positive constant that incorporates the main sources of error in the quantum hardware, so that all atoms are ensured to start in the ground state |$\ket{0}$. More importantly, the final detuning should lie in a narrow interval $[\Delta_{\text{LB}},\Delta_{\text{UB}}]$, where the bounds are explicitly given by
$\frac{C_{6}}{12\,a^{6}}\leq\Delta_{\text{f}}\leq\frac{C_{6}}{8\,a^{6}}$ (3)
in the case of an underlying square lattice with a distance $a$ between adjacent nodes, see Appendix A. Note that such bounds are not universal, but they are satisfied by the vast majority of unit disk graphs for a given lattice.
@@
A detailed analysis of the classical dynamics of the simulated annealing for MIS reveals that the graphs with many suboptimal solutions are likely to trap the algorithm in local minima, hence missing the global solution [19]. This observation leads to the following definition of the hardness parameter of a graph:
@@
2 Numerical optimization
We consider a fixed set of 500 representative graphs of size 8 to 17, which allows us to perform fast numerical simulations and to use efficient deterministic optimization algorithms, see Appendix C. This dataset is such that the correlation between the size of the graphs and their hardness parameter is minimized. Therefore, the results that depend principally on $\mathcal{HP}$ can be extrapolated, to a certain extent, to graphs of any size.
In this work, we focus on the probability to reach a MIS solution at the end of the adiabatic evolution, which we denote by $P_{\text{MIS}}$ and use as the score function to be optimized for some parametrization $\Theta$ of the schedules:
$P_{\text{MIS}}(\Theta)=\sum_{i=1}^{D_{|\text{MIS}|}}\big{|}\,\braket{\psi(\Theta,t_{\text{f}})\,}{\,\text{MIS}_{i}}\,\big{|}^{2},$ (5)
where |$\ket{\psi(\Theta,t_{\text{f}})}$ is the final quantum state and |$\ket{\text{MIS}_{i}}$ is the basis vector that corresponds to the $i$-th solution of the MIS problem. Note that most studies consider the approximation ratio instead (normalized mean energy of the final state), but in our opinion only the lowest energy level and not the full energy spectrum has a practical utility because of the very definition of the hardness parameter in Eq. (4).
Following the “best practices” listed in [21], we consider protocols for which the Rabi amplitude $\Omega(t)$ reaches $\Omega_{\max}$ during the evolution, so that the associated dynamic blockade radius becomes $R_{\text{b}}=(C_{6}/\Omega_{\max})^{1/6}\approx 8.4$. In order to generate a unit disk graph on the nearest and next-nearest neighbors of a square lattice, this radius should satisfy $\sqrt{2}a<R_{\text{b}}<2a$, and consequently $5.94.2<a<$ 5.9. We set $a=5$, which is close to the geometric mean of the bounds and which yields good results in preliminary calculations. We consider a fixed evolution time $t_{\text{f}}=1$ so that $P_{\text{MIS}}$ varies as much as possible depending on the various protocols; much smaller or larger times lead to either extremely bad or extremely good results independently of the precise shape of the schedules.
@@ 2 Numerical optimization
2.1 Piecewise linear protocol
The typical adiabatic protocol is based on piecewise linear schedules, where $\Omega$ is slowly ramped up, then $\Delta$ is ramped from negative to positive, and finally $\Omega$ is ramped down, see [21, Fig. 3.1b] or [35, Fig. 2c]. Usually, these schedules are symmetric in the sense that the initial and final ramp times $\tau_{\text{i}}$ and $\tau_{\text{f}}$ are equal, and that the initial and final detunings $\Delta_{\text{i}}$ and $\Delta_{\text{f}}$ have opposite values.
In our case, we are looking for schedules that are as general as possible, so we keep all parameters independent. Moreover, we consider distinct values of the ramp times for $\Omega$ and $\Delta$. After few optimization steps, it appears that the ramp times of the detuning tend to vanish; we therefore set them to zero. Then, we optimize the four remaining parameters $(\tau_{\text{i}},\tau_{\text{f}},\Delta_{\text{i}},\Delta_{\text{f}})$ for each graph of the dataset. A noticeable observation is that the optimum parameters depend mostly on the hardness parameter and not on the size of the graph, see Fig. 1. The consequence is twofold: first, we can fit these data to generate a model of optimized schedules, as illustrated in Fig. 2. Second, such a model derived from small graphs can be extrapolated to larger graphs, see Sec. 3.2.
@@ 2.1 Piecewise linear protocol
Figure 2: Optimized piecewise linear schedules with 4 free variables. The Rabi amplitude (top) and the detuning (bottom) depend on the hardness of the graph and are illustrated here for $\mathcal{HP}\in\{0.5,1,2,4,8\}$ (thin lines, from green to maroon) and in the limit $\mathcal{HP}\to\infty$ (dark maroon thick line). The optimized values become in this limit $\tau_{\text{i}}\approx t_{\text{f}}/2$, $\tau_{\text{f}}\approx t_{\text{f}}/10$, $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/2$, and $\Delta_{\text{f}}=\Delta_{\text{UB}}$.
Figure 3: Optimized piecewise linear schedules with 6 free variables. In this case, $\tau_{\text{i}}$ dislays a clear dependence on $\mathcal{HP}$. In the limit of infinite hardness, the optimized values are $\tau_{\text{i}}\approx 0.22\,t_{\text{f}}$ and $\tau_{\text{f}}\approx 0.12\,t_{\text{f}}$ for the Rabi amplitude (top) and $\Delta_{\text{i}}\approx-0.83\,\Delta_{\text{UB}}$, $\tau_{\text{m}}\approx 0.34\,t_{\text{f}}$, $\Delta_{\text{m}}\approx 0.36\,\Delta_{\text{UB}}$, and $\Delta_{\text{f}}=\Delta_{\text{UB}}$ for the detuning (bottom).
The hardness-dependent linear schedules yield very good results compared to existing protocols, which motivates us to go one step further by considering two more degrees of freedom $(\tau_{\text{m}},\Delta_{\text{m}})$ for the detuning, see Fig. 3. The resulting optimized schedules are still easily implementable, yet they perform much better than all benchmark protocols, see Sec. 2.3.
@@
The aim of the adiabatic evolution is to find the MIS of a graph, which in turn allows one to calculate the hardness of that very graph. In the case of small graphs, since all vertex configurations are enumerated effortlessly, we consider that $\mathcal{HP}$ is a known value that can be used to choose the best schedules from the fitted model. For large graphs, however, the hardness parameter cannot be considered to be known beforehand. In that case, one should either take the schedules in the limit $\mathcal{HP}\to\infty$, which yields quite good results for all graphs, or eventually apply a one-dimensional search on the hardness parameter, which is still much easier than performing a multi-dimensional optimization of all schedule parameters.
2.2 Counterdiabatic protocol
@@
The limited coherence time of the quantum state in experiments is a major source of error in adiabatic computing. Well-established methods to overcome such limitations include the shortcuts to adiabadicity, among which is the counterdiabatic (CD) driving [36]. The idea of the CD driving is to add a gauge potential to the adiabatic Hamiltonian such that the transitions between eigenstates are suppressed during the evolution. However, the calculation of the exact gauge potential is a very difficult task in general, and these mathematical terms may not be always adapted to an experimental implementation. Nevertheless, by neglecting the interactions of the Rydberg Hamiltonian, one is able to find a suitable gauge potential which utilizes the phase of the Rabi drive [28]. In this study, we propose to generalize the adiabatic schedules of this so-called Analog Counterdiabatic Quantum Computing (ACQC) protocol as follows:
$\Omega_{\text{ad}}(t)$ $=\Omega_{\max}\,\sin\left(\frac{\pi}{2}\,\sin\left(\theta_{\Omega}(t)\right)\right)^{2},$ (6a)
$\Delta_{\text{ad}}(t)$ $=\frac{\Delta_{\text{i}}+\Delta_{\text{f}}}{2}+\frac{\Delta_{\text{i}}-\Delta_{\text{f}}}{2}\cos\left(\theta_{\Delta}(t)\right),$ (6b)
where $\theta(t)$ is any monotonic smooth function that ranges from 00 to $\pi$; here we set $\theta_{\Omega,\Delta}(t)=\pi\,t/t_{\text{f}}$. Moreover, based on the variational procedure described in [27], we improve on the ACQC protocol by including the Rydberg interactions in the calculation of the CD terms, thus adapting the gauge potential to each different graph, see Appendix B. The detuning is not altered by this operation, whereas the Rabi drive becomes:
$\Omega$ $=\sqrt{\Omega_{\text{ad}}^{2}+\Omega_{\text{cd}}^{2}},$ (7a)
$\phi$ $=\arctan(\Omega_{\text{cd}}/\Omega_{\text{ad}}),$ (7b)
with
$\Omega_{\text{cd}}=\frac{\Omega_{\text{ad}}\dot{\Delta}_{\text{ad}}-\dot{\Omega}_{\text{ad}}\Delta_{\text{ad}}+\dot{\Omega}_{\text{ad}}\,T_{1}\,\nu}{\Omega_{\text{ad}}^{2}+\Delta_{\text{ad}}^{2}-2\Delta_{\text{ad}}\,T_{1}\,\nu+T_{2}\,\nu^{2}}$ (8)
and where $0\leq\nu\leq C_{6}/a^{6}$ is an additional free parameter that interpolates between the standard and the generalized gauge potentials. In the latter equation, the variables $T_{1}$ and $T_{2}$ depend on the interaction terms $V_{ij}$ and are therefore specific to the graph to be solved. As expected, the CD contribution vanishes for $t_{\text{f}}\to\infty$ since $\dot{\Omega}_{\text{ad}}$ and $\dot{\Delta}_{\text{ad}}$ tend to zero in this limit.
Similarly as for the piecewise linear protocol, we optimize the schedule parameters $(\Delta_{\text{i}},\Delta_{\text{f}},\nu)$ for all graphs of the dataset, see Fig. 4. The interpolating coefficient $\nu$ tends to a strictly positive value for the hardest graphs, indicating that the generalized gauge potential plays an important role in the CD protocol. Note that the easy graphs exhibit optimized values of $\Delta_{\text{f}}$ that are smaller than the statistically determined constant $\Delta_{\text{LB}}$. This is due to the fact that the exact lower bounds on the detuning are much smaller in that case, see Fig. 11 in the appendix. The shape of the CD phase is highly nontrivial, and we observe a strong shift toward higher values of the detuning for hard MIS instances, see Fig. 5. Finally, we also tried to further improve the results by considering more evolved $\theta$ functions—in particular, to mimic the skewed shape of the Rabi amplitude in Fig. 2—but no significant gain was found in this case.
@@
In Aquila, the Rabi amplitude and the detuning must be defined as piecewise linear and not smooth functions, while the phase must be piecewise constant. However, the step size can be as small as 0.05 µs, which results in 80 intervals for the longest coherent evolution. The schedules are therefore well approximated in the quantum hardware, but one could slightly improve the experimental results by taking into account this hardware capability in the numerical simulations.
@@
Figure 4: Optimum parameters for the CD schedules. The initial detuning (top) reaches the upper bound $-\delta_{\Delta}$ for several graphs but tends to a lower value $\Delta_{\text{i}}\approx-\Delta_{\text{UB}}/10$ for the hardest graphs. The final detuning (bottom) also increases with the hardness parameter and quickly reaches the limit $\Delta_{\text{UB}}$. The coefficient $\nu$ (not shown) displays a similar behavior and tends to $\nu=4.30$ in the current setting.
Figure 5: Optimized counterdiabatic schedules with 3 free variables. Unlike the detuning (bottom), the Rabi amplitude (top) and phase (middle) depend not only on $\mathcal{HP}$ but also on $t_{\text{f}}$ and on the graph-dependent variables $T_{1}$ and $T_{2}$. Here, the curves are displayed for $t_{\text{f}}=1$ and for the mean values of the dataset $\overline{T}_{1}\approx 0.93$ and $\overline{T}_{2}\approx 1.46$. The light green area represents the variation due solely to $T_{1}$ and $T_{2}$ for $\mathcal{HP}=0.5$.
2.3 Results